OpenStudy (anonymous):

Please Help ??? Solve the following triangle. A=28 degrees B=20 degrees c= 17 m FIND Angle C

rebeccaxhawaii (rebeccaxhawaii):

what is 17m?

OpenStudy (anonymous):

side C

OpenStudy (dinamix):

c= (180-(a+b)) the angle c = 180-(20+28) @hamza123

OpenStudy (anonymous):

132

OpenStudy (anonymous):

what about side a

OpenStudy (mrnood):

you can find a using the sin rule a/sinA = c/sinC

OpenStudy (anonymous):

a/sin28=17m/sin132

OpenStudy (dinamix):

@MrNood i was putt my draw but u are fast cool

OpenStudy (anonymous):

so what do i do next

OpenStudy (mrnood):

you can solve all the unknowns from the sin rule a/sin28=17m/sin132 put in the values for sin 28 and sin 132 and oyu can get a from your equation

OpenStudy (mrnood):

17m I assume is 17 metres - so you have the only unknown is a

OpenStudy (anonymous):

i cant find the answer because i dont know how to set it up

OpenStudy (mrnood):

you wrote the equation: a/sin28=17m/sin132 multiply both sides by sin 28

OpenStudy (anonymous):

im still confused

OpenStudy (anonymous):

so multiply 17m/sin132 * sin 28

OpenStudy (anonymous):

86.75 is the answer?

OpenStudy (anonymous):

is this correct

OpenStudy (anonymous):

17 *sin132*sin28

OpenStudy (anonymous):

should the answer be 0.2445

OpenStudy (anonymous):

what is there no solution??? pls answer

OpenStudy (mrnood):

you have written the correct equation - you have done all the 'triangle bits' a/sin28=17m/sin132 you simply have to solve your equation for a \[\frac{ a }{\sin28 } = \frac{17 }{ \sin132 }\] solve it like any equation - you should be able to do these equations if you have already reached the trigonometry stage in maths

OpenStudy (jdoe0001):

hmmm

OpenStudy (mrnood):

I gave the hint above: multiply both sides by sin 28

OpenStudy (jdoe0001):



OpenStudy (jdoe0001):

for, if that's that, then you can simply use the internal angles sum

OpenStudy (anonymous):

sin 132a=17sin28 then solve

OpenStudy (mrnood):

@jdoe0001 that has already been done - the angle is 132 the sin rule is all that is needed but she has even done THAT all she needs is to solve the equation for a

OpenStudy (jdoe0001):

ohh ahemm... alrite =)

OpenStudy (jdoe0001):

the posting only says, find C

OpenStudy (anonymous):

no a

OpenStudy (anonymous):

so 132=17sin128

OpenStudy (mrnood):

just multiply the equation YOU wrote on both sides by sin 28 take it steady and be methodical (your last post has missed out the a)

OpenStudy (anonymous):

so 132 *sin 28

OpenStudy (mrnood):

\[\frac{ a }{ \sin 28} = \frac{ 17 }{ \sin 132 }\] multiply both sides by sin 28 BE METHODICAL _ DONT GUESS

OpenStudy (anonymous):

17sin28/sin132 sin28

OpenStudy (mrnood):

no I'm sorry - you are studying trigonometry - this is way beyond doing simple crossmultiplying look at the equation - wite *sin 28 on EACH SIDE then simplify

OpenStudy (anonymous):

for both 17 and sin 132

OpenStudy (mrnood):

you have an equation with a in it as the only unknown to solve it you want an equation that reads a=???? to get there take the equation I werote in my last post and write "times sin 28" on EACH SIDE

OpenStudy (mrnood):

as an example a/2 = 17/3 multiply both sides by 2 a*2/2 = 17*2/3 the left side cancels so a = 34/3 = 11.33333 do the same with sin 28 in your equation